# Machine Learning Symmetry Discovery for Classical Mechanics

### v1

@@ III.1 Kepler problem
Following the pipeline outlined in Sec. II, we tested multiple conjectured symmetry group dimensions, $n=6$, $n=10$, and $n=15$  [^2] An $\mathrm{SO}(3)$ symmetry with $n=3$ is obvious since the Kepler Hamiltonian preserves the angular momentum with rotation symmetry., corresponding to $\mathrm{SO}(4)$, $\mathrm{SO}(5)$, and $\mathrm{SO}(6)$, respectively. We then trained the neural networks on the simulation dataset using the loss function in (11) with $\alpha=1$ and $\beta=1\times 10^{-4},2\times 10^{-4}...,10^{-3}$ as shown in Fig. 3.
Figure 2: Examples of time evolution driven by the learned Hamiltonian. Each sub-figure compares the time evolution generated by the true Hamiltonian $H_{\text{Kep}}$ (dashed gray) and the learned Hamiltonian $G_{0,\theta}$ (blue), both starting from the same initial condition ${\bm{x}}(t=0)$. The evolution trajectories are plotted after projecting the $3d$ variables onto a $2d$ plane.
A straight forward method to evaluate the Hamiltonian learning is by reproducing the time evolution trajectories, as illustrated in Fig. 2. Starting from the same initial condition ${\bm{x}}(t=0)$, the well-trained $G_{0,\theta}$ is able to reproduce the time evolution Eq. (6) highly identical to that of $H$ for a finite amount of time, before losing track due to potentially accumulating errors.
By tracking the converged values of each term in Eq. (11) for $n=6$, $n=10$, and $n=15$, we find that only the $n=6$ conjecture with $\beta=1\times 10^{-4},2\times 10^{-4},3\times 10^{-4}$ can minimize $\mathcal{L}_{\text{G}}$ to zero while keeping $\mathcal{L}_{\text{I}}$ finite, as shown in Fig. 3. This suggests a potential symmetry group of dimension $n=6$ that is worth further investigation.
Figure 3: The sub-figures (a), (b), and (c) show $\mathcal{L}_{\text{G}}$ versus training epochs for $n=6$, $n=10$, and $n=15$. (d) shows the converged $\mathcal{L}_{\text{G}}$ versus $\beta$. The $\mathcal{L}_{\text{G}}(n=6)$ with $\beta=1\times 10^{-4},2\times 10^{-4},3\times 10^{-4}$ are able to converge to near zero comparing to other trials.
To further identify the learned symmetry group, we diagonalize the Killing form matrix $\mathbf{B}=\mathbf{U}\mathbf{D}\mathbf{U}^{T}$, whose entries defined as $B^{il}=\sum_{k,j}f^{ijk}f^{lkj}$, where $f^{ijk}$ is the machine-learned structure coefficient in Eq. (2). Example results yields $\mathbf{D}=\text{diag}(-3.40,-3.44,-3.47)\oplus\text{diag}(-1.22,-1.20,-1.17)$. This pattern is consistent with theoretical expectations of the decoupling of the two $\mathfrak{su}(2)$ algebras, as $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$. Further projecting the structure coefficient using $\mathbf{U}$ reproduces the $\mathfrak{so}(4)$ algebra as shown in Fig. 4. A detailed explanation of the method used to identify this symmetry group is provided in Appendix B.
@@ III.2 Harmonic oscillator
We apply the same method to test multiple conjectured symmetry group dimensions $n=8$, $n=10$, $n=15$. The results of Hamiltonian learning are presented in Fig. 5. The learning task is notably easier in the harmonic oscillator system compared to the Kepler problem, as the Hamiltonian and conserved quantities in the harmonic oscillator have a simpler quadratic form in terms of ${\bm{p}}$ and ${\bm{q}}$.
At $n=8$, the 8 eigenvalues of the Killing form of the structure coefficients exhibit strong degeneracy, for example: $\mathbf{D}=\mathrm{diag}$($-18.48$, $-18.46$, $-18.44$, $-18.43$, $-18.43$, $-18.41$, $-18.40$, $-18.38$). The structure coefficient reconstructs the $\mathfrak{su}(3)$ Lie algebra as shown in the Fig. 6. This strongly indicates a hidden $\mathrm{SU}(3)$ symmetry in the harmonic oscillator system.
Figure 5: Examples of time evolution driven by the learned Hamiltonian in harmonic oscillator system.
@@
Figure 7: The sub-figures (a), (b), and (c) show $\mathcal{L}_{\text{G}}$ versus training epochs for $n=8$, $n=10$, and $n=15$. (d) shows the converged $\mathcal{L}_{\text{G}}$ versus $\beta$. The $\mathcal{L}_{\text{G}}(n=8)$ with $\beta=1\times 10^{-4},2\times 10^{-4},3\times 10^{-4},,4\times 10^{-4}$ are able to converge to near zero comparing to other trials.
@@ IV Summary and Discussion
  • Symmetry discovery without prior knowledge
    In this study, we propose a data-driven, deep learning-based Machine Learning Symmetry Discovery algorithm designed to explore and analyze continuous symmetries in classical mechanical systems. Previous works[12, 13, 14, 11, 15, 16, 21, 22] have demonstrated methods to verify conjectured continuous symmetries of a given Hamiltonian. To make further improvements, our pipeline requires only the conjectured dimension of the symmetry group. The machine learning algorithm can then discover the symmetry transformations and subsequently identify the symmetry group, whether Abelian or non-Abelian, by analyzing the Killing form of the structure coefficients. This study addresses a crucial topic: discovering hidden symmetries within data itself using machine learning. Many continuous symmetries in physical systems are often obscure and difficult to uncover, even when the explicit form of the Hamiltonian is known. By leveraging deep learning, we aim to surpass human capabilities in identifying these symmetries. Our approach enables the algorithm to learn directly from the data, uncovering potential underlying symmetries that might otherwise remain hidden through traditional analysis. This represents a significant advancement in the application of machine learning to the discovery of complex patterns in physical systems.
  • Extension to quantum and many-body system
@@
  • Limitations
    Many symmetry transformations driven by quantitys have highly nonlinear dependence on the canonical coordinates. If the target symmetry transformation exceeds the representational capacity of linear feed-forward neural networks, then more advanced neural networks [7, 26, 27, 28, 29] are required for symmetry discovery. Otherwise, the learning process may be hindered. Developing neural networks specifically tailored for AI-driven scientific tasks is crucial, as these tasks often involve complex structures and patterns that demand greater representational power and precision compared to general-purpose neural networks.
    The MLSD algorithm relies on an independence loss term $\mathcal{L}_{\text{I}}$ to encourage the search for linearly independent symmetry transformations. This independence loss plays a crucial role in guiding the model to discover distinct symmetries. However, despite being weighted by a small parameter $\beta$, this term can still introduce instability into the gradient descent process, making the optimization unstable and potentially causing the task to fail. A better design for the symmetry searching task is still worth exploring.
@@ Appendix A Symmetry Analysis
    These relations demonstrate that ${\bm{\mathcal{L}}}$ and ${\bm{\mathcal{R}}}$ each satisfy the commutation relations of the $\mathfrak{su}(2)$. Consequently, we have two decoupled $\mathfrak{su}(2)$, where
@@ Appendix B Basis transformation of Lie algebra
    There are two methods to identify the Lie group after training with MLSD. Suppose $f$ is the structure coefficient obtained from MLSD and $\epsilon$ represents a reference coefficient. The straightforward approach is to use gradient descent to find an optimal transformation matrix $M$ by minimizing the objective function: $\sum_{a,b,c}\|(M^{-1})_{a}^{i}(M^{-1})_{b}^{j}f_{ij}^{k}M_{k}^{c}-\epsilon_{ab}^{c}\|$. If such a matrix $M$ exists, than the two groups are isomorphism to each other.
@@
    to obtain $\mathbf{D}$. The eigenvalues of the Killing form matrix $\mathbf{B}$ will exhibit degeneracy corresponding to the sub-algebra structure of ${\bm{G}}$. For instance, in the Kepler problem, the symmetry algebra is $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$ Consequently, diagonalizing the Killing form matrix of the trained structure coefficients $f$ should yield two sets of degenerate eigenvalues: three identical eigenvalues corresponding to one $\mathfrak{su}(2)$ sub-algebra and another three identical eigenvalues corresponding to the other $\mathfrak{su}(2)$ sub-algebra. This eigenvalue pattern reflects the direct sum structure of the algebra. In the case of the harmonic oscillator problem, the symmetry algebra is $\mathfrak{su}(3)$ , which cannot be decomposed into a direct sum of sub-algebras. Therefore, the eigenvalues of the Killing form matrix should exhibit a single set of eight identical values, reflecting the irreducibility of the $\mathfrak{su}(3)$ algebra.

### v2

@@ III.1 Kepler problem
Following the pipeline outlined in Sec. II, we tested multiple conjectured symmetry dimensions $n=4,6,8,10,$ and $12$. For each choice of $n$, the neural networks were trained on the same simulation dataset using the loss function in (13) as shown in Fig. 3. For each $n$, we repeat training with $15$ random seeds; error bars indicate the standard deviation across seeds. Among these, $n=6$ corresponds to the expected $\mathrm{SO}(4)$ symmetry of the Kepler problem, while larger values such as $n=10$ represent over-parameterized hypotheses. The remaining cases ($n=4,8,12$) do not correspond to standard symmetry algebras and are included solely as complementary tests.
Figure 2: Hamiltonian learning of the 3D Kepler problem: orbit comparison. Two representative examples (a, b) showing ground truth orbits (gray) versus learned trajectories (blue) integrated over one orbital period $T$. The red star marks the central body at the origin. Blue arrowheads indicate the velocity direction at the end of integration. Right panels show the corresponding energy drift $|\Delta H|$ as a function of normalized time $t/T$, demonstrating that errors accumulate due to imperfect Hamiltonian learning. All trajectories are integrated using the velocity-Verlet symplectic integrator.
A straightforward way to assess the Hamiltonian learning is by comparing the resulting time-evolution trajectories, as illustrated in Fig. 2. Starting from the same initial condition ${\bm{x}}(t=0)$, the learned Hamiltonian $G_{0,\theta}$ reproduces the local time evolution generated by the true Hamiltonian $H$ with good agreement over short to intermediate times. Small deviations become apparent at longer times, which can be attributed to optimization inaccuracies in the learned Hamiltonian $G_{0,\theta}$ together with the accumulation of local numerical errors, even when using a symplectic integrator. While these effects influence long-time trajectory accuracy, they remain tolerable for the present tasks and do not obstruct the identification of the underlying symmetry structure.
Tracking the converged loss values in Eq. (13) reveals a minimum at $n=6$ among the tested dimensions $n=4,6,8,10,$ and $12$ (Fig. 3). While $\mathcal{L}_{\text{G}}$ does not vanish exactly due to optimization limitations, the marked separation between $n=6$ and other cases indicates that the correct symmetry dimension is $n=6$.
Figure 3: MLSD for the 3D Kepler problem: training convergence and symmetry dimension identification. (a–e) Training loss $\mathcal{L}_{G}$ versus epoch for symmetry dimensions $n=4,6,8,10,12$, with 15 independent random seeds per dimension. Insets show the distribution of final converged losses. (f) Mean converged loss versus symmetry dimension $n$, with error bars indicating standard deviation across seeds. The vertical dashed red line marks the true symmetry dimension $n=6$, which achieves the lowest loss, correctly identifying the $\mathfrak{so}(4)$ symmetry of the 3D Kepler problem.
To further identify the learned symmetry group, we analyze the machine-learned structure coefficients $f^{ijk}$ using two complementary post-processing procedures described in Appendix B: (i) diagonalization of the Killing form and (ii) optimization of a linear transformation that maps the learned coefficients to a standard Lie-algebra basis. In practice, both approaches yield indistinguishable identification of the underlying algebra for the problems studied here.
Specifically, we diagonalize the Killing form matrix $\mathbf{B}=\mathbf{U}\mathbf{D}\mathbf{U}^{T}$, whose entries are defined as $B^{il}=\sum_{k,j}f^{ijk}f^{lkj}$. A representative result yields $\mathbf{D}=\mathrm{diag}(-3.40,-3.44,-3.47)\oplus\mathrm{diag}(-1.22,-1.20,-1.17)$, which is consistent with the expected decoupling of two $\mathfrak{su}(2)$ subalgebras, since $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$. Further projecting the structure coefficients using $\mathbf{U}$ reproduces the standard $\mathfrak{so}(4)$ algebra, as shown in Fig. 4. For clarity, we present the Killing-form-based identification here, while both methods were applied for all reported results.
@@ III.2 Harmonic oscillator
We apply the same procedure to the harmonic oscillator, testing symmetry dimensions $n=6,7,8,9,$ and $10$. For each $n$, we repeat training with $15$ random seeds; error bars indicate the standard deviation across seeds. The learned Hamiltonian dynamics (Fig. 5) show smaller long-time deviations than in the Kepler case due to the simpler quadratic form of the harmonic oscillator Hamiltonian.
At $n=8$, the 8 eigenvalues of the Killing form of the structure coefficients exhibit strong degeneracy, for example: $\mathbf{D}=\mathrm{diag}$(18.48-18.48, 18.46-18.46, 18.44-18.44, 18.43-18.43, 18.43-18.43, 18.41-18.41, 18.40-18.40, 18.38-18.38). The structure coefficient reconstructs the $\mathfrak{su}(3)$ Lie algebra as shown in the Fig. 6. This strongly indicates a hidden $\mathrm{SU}(3)$ symmetry in the harmonic oscillator system.
Figure 5: Hamiltonian learning of the 3D harmonic oscillator: orbit comparison. Two representative examples (a, b) showing ground truth orbits (gray) versus learned trajectories (blue) integrated over one period $T$. The red star marks the equilibrium position at the origin. Blue arrowheads indicate the velocity direction at the end of integration. Right panels show the corresponding energy drift $|\Delta H|$ as a function of normalized time $t/T$, demonstrating error accumulation from imperfect Hamiltonian learning. All trajectories are integrated using the velocity-Verlet symplectic integrator.
@@
Figure 7: MLSD for the 3D harmonic oscillator: training convergence and symmetry dimension identification. (a–e) Training loss $\mathcal{L}_{G}$ versus epoch for symmetry dimensions $n=6,7,8,9,10$, with 15 independent random seeds per dimension. Insets show the distribution of final converged losses. (f) Mean converged loss versus symmetry dimension $n$, with error bars indicating standard deviation across seeds. The vertical dashed red line marks the true symmetry dimension $n=8$, which achieves the lowest loss, correctly identifying the $\mathfrak{su}(3)$ symmetry of the 3D harmonic oscillator.
@@ IV Summary and Discussion
IV.1 Symmetry discovery without prior knowledge
In this study, we propose a data-driven, deep learning-based Machine Learning Symmetry Discovery algorithm designed to explore and analyze continuous symmetries in classical mechanical systems with integrable Hamiltonian dynamics. Previous works[14, 13, 23, 22, 17, 4, 5, 24] have demonstrated methods to verify conjectured continuous symmetries of a given Hamiltonian. Within this setting, our pipeline requires only the conjectured dimension of the symmetry group as input. The machine learning algorithm can then discover candidate symmetry transformations and subsequently identify the associated symmetry algebra, whether Abelian or non-Abelian, by analyzing the Killing form of the learned structure coefficients.
In a controlled benchmark setting, MLSD discovers candidate conserved quantities from trajectory data and recovers the associated Lie-algebraic structure via learned structure coefficients and Killing-form analysis. The contribution is a practical end-to-end pipeline, including reproducibility diagnostics across symmetry-dimension hypotheses, that enables systematic identification of non-Abelian symmetry algebras from data in integrable Hamiltonian systems.
IV.2 Choice of examples and limitations
The present implementation of MLSD is constrained by the expressive capacity of the neural networks used to parametrize $G_{i}({\bm{x}})$.
The Kepler problem and the three-dimensional harmonic oscillator were chosen as demonstrations because they provide canonical benchmark systems and analytically well-understood symmetry structure, making them suitable test cases for validating the proposed framework. In systems with more complex or less regular dynamics, including mixed or chaotic phase space, the relevant invariants may involve higher-order nonlinearities, non-polynomial structure, or regime-dependent effective descriptions.
If the target symmetry transformations or the Hamiltonian $H({\bm{x}})$ exceed the representational capacity of the chosen neural-network hypothesis class, the current implementation of MLSD may fail to recover the correct symmetry structure.
All results in this work are obtained using clean simulated trajectories expressed in canonical coordinates. The robustness of MLSD to measurement noise, partial observability, or imperfect coordinate reconstruction is not addressed here. Understanding how approximate symmetries degrade under noise and identifying noise thresholds beyond which symmetry discovery fails are important and nontrivial directions for future work.
The MLSD algorithm also relies on an independence loss term $\mathcal{L}_{\text{I}}$ to encourage the discovery of linearly independent symmetry generators. Although weighted by a small parameter $\beta$, this term may introduce optimization sensitivity and can affect convergence behavior. In addition, symmetry-dimension selection in the present work is based on statistical comparison of converged loss values across multiple random initializations rather than formal hypothesis testing. While this provides a reproducible and practically effective criterion, developing more rigorous model-selection or uncertainty-quantification frameworks is an important avenue for future investigation.
Finally, the coverage of the training dataset plays a critical role in determining the symmetry structure recovered by MLSD. When trajectories sample diverse regions of phase space within a regular dynamical regime, the method is encouraged to recover symmetries of the global Hamiltonian. In contrast, for datasets restricted to narrow regions of phase space—including trajectories confined to KAM islands or chaotic subsets—the method may instead infer an effective, local symmetry algebra valid only within that region. Addressing these challenges through improved data-sampling strategies, more stable optimization objectives, and more expressive architectures represents an important direction for future work.
IV.3 Extension to quantum and many-body system
@@ Appendix A Symmetry Analysis
[…]
@@ Appendix B Basis transformation of Lie algebra
    There are two methods to identify the Lie group after training with MLSD. Suppose $f$ is the structure coefficient obtained from MLSD and $\epsilon$ represents a reference coefficient. The straightforward approach is to use gradient descent to find an optimal transformation matrix $M$ by minimizing the objective function: $\sum_{a,b,c}\|(M^{-1})_{a}^{i}(M^{-1})_{b}^{j}f_{ij}^{k}M_{k}^{c}-\epsilon_{ab}^{c}\|$. If such a matrix $M$ exists, then the two groups are isomorphism to each other.
@@
    to obtain $\mathbf{D}$. The eigenvalues of the Killing form matrix $\mathbf{B}$ will exhibit degeneracy corresponding to the sub-algebra structure of $\bm{G}$}. For instance, in the Kepler problem, the symmetry algebra is $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$ . Consequently, diagonalizing the Killing form matrix of the trained structure coefficients $f$ should yield two sets of degenerate eigenvalues: three identical eigenvalues corresponding to one $\mathfrak{su}(2)$ sub-algebra and another three identical eigenvalues corresponding to the other $\mathfrak{su}(2)$ sub-algebra. This eigenvalue pattern reflects the direct sum structure of the algebra. In the case of the harmonic oscillator problem, the symmetry algebra is $\mathfrak{su}(3)$ , which cannot be decomposed into a direct sum of sub-algebras. Therefore, the eigenvalues of the Killing form matrix should exhibit a single set of eight identical values, reflecting the irreducibility of the $\mathfrak{su}(3)$ algebra.